# Online Decision-Focused Learning

### v1

@@ Appendix D A stronger version of H 2 for a stronger notion of regret
We know by Proposition 2 that the regret of Algorithm 1 satisfies the following asymptotic bound:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{1+P_{T}}{\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+n\sum_{t\in[T]}\alpha_{t}+\xi T}\right]}\right)\;.$
The rest of the proof proceeds exactly as the one of Corollary 1. ∎
F.3 Proofs of Appendix B.
See 4
Proof.
For the sake of conciseness we use the noation $\mathbb{E}_{t}$ to denote $\mathbb{E}\mathopen{}\mathclose{{}\left[\cdot\mid\mathcal{H}_{t-1}}\right]$ We have, for any $\theta$, the following decomposition of the static regret
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle}\right]$ (47)
@@
Then, taking the $\sup$ over $\Theta$ in Equations 47 and 49, and defining
@@
in Equation 48 yields for any $\theta$:
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle}\right]$
Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$
Since for any $t\in[T]$, $\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{2}\leqslant D_{\mathcal{Z}}$ by H 1-(ii),
$\mathfrak{R}_{T}^{\mathrm{S}}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}\;$
First, notice that, by Lemma 3, and because for any $t,\alpha_{t}=\alpha$ we have for all $t$:
$\mathbb{E}_{X_{1},\cdots,X_{T}}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$ $\leqslant 2D_{\mathcal{Z}}\alpha\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}\right)}\right)\;.$ (50)
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}^{\mathrm{S}}_{T}(\theta)}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}^{\mathrm{S}}_{T}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}\right)}\right).$ (51)
To bound the first term in the right-hand-side of Equation 51. by the definition of conditional expectation, we have:
$\mathrm{Reg}^{\mathrm{S}}_{T}$ $=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}\right\rangle-\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle}\right]$
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}^{\mathrm{S}}_{T}}\right]$ $\leqslant\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{}\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}\right].$
@@
One recognises (up to a factor $T$) the left-hand side of Proposition 3 on the loss sequence $(\tilde{f}_{t})_{t\in[T]}$. Furthermore, we can use this proposition as, given our choice of $\mathcal{R}$, for any $t\in[T]$, $\tilde{f}_{t}$ is $L-$Lipschitz with $L=5D_{\mathcal{Z}}G(4\alpha)^{-1}$ almost surely by Lemma 2. We then have:
@@
$\mathbb{E}_{u_{1},\cdots,u_{T}}\mathopen{}\mathclose{{}\left[\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}\right]\leqslant\mathcal{O}\mathopen{}\mathclose{{}\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi m\frac{1}{\alpha}}\right)\;.$ (52)
Dividing Equation 51 by $T$, and plugging Equation 52 into it concludes the proof. ∎
F.4 Proofs of Appendix D.
Proof of Lemma 1.
The proof starts exactly in the same way as the one of Lemma 3, where proved in (24) that:
Moreover, we have for any $\varepsilon>0$:
$1-C_{0}\varepsilon^{\beta}$ $\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\;\inf_{j\neq I_{t}}u_{j}(\theta,X_{t})-u_{I_{t}}(\theta,X_{t})\geqslant\varepsilon\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
$\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\;\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}{\mathcal{H}_{t-1}}\leqslant 2\alpha_{t}\ln(d)\varepsilon^{-1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
$\leqslant\mathbb{P}\mathopen{}\mathclose{{}\left(\;\sup_{\theta\in\Theta}\;\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}\leqslant 2\alpha_{t}\ln(d)\varepsilon^{-1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right)$
where the first inequality stems from H 4, and the second one from Lemma 4. It follows that:
$\mathbb{P}\mathopen{}\mathclose{{}\left(\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}>2\frac{\alpha_{t}\ln(d)}{\varepsilon}\mid\mathcal{H}_{t-1}}\right)$ $=1-\mathbb{P}\mathopen{}\mathclose{{}\left(\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\leqslant 2\frac{\alpha_{t}\ln(d)}{\varepsilon}\mid\mathcal{H}_{t-1}}\right)$
$\leqslant C_{0}\varepsilon^{\beta}\;,$ (54)
Hence, plugging (54) in (24) gives:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)\|_{1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right]$ $\leqslant\alpha_{t}+2\alpha_{t}\ln(d)C_{0}\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon\;.$
And the result is established by bounding $\varphi(\beta)=\int_{\alpha_{t}\ln(d)}^{2\ln(d)}\varepsilon^{\beta-2}\,\mathrm{d}\varepsilon$ in the same way as in the proof of Lemma 3. ∎
Proof of Proposition 5.
Observe that the pseudo dynamic regret can be decomposed as follows:
$\mathfrak{R}_{T}^{\mathrm{d}}$ $=\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle$ (55)
$+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}\right\rangle-\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle$ (56)
$+\sum_{t\in[T]}\mathopen{}\mathclose{{}\left[\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}\right\rangle-\inf_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)}\right\rangle}\right]$ (57)
Taking the $\sup$ over $\Theta$ and defining by ~$\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}:=\sum_{t\in[T]}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle$ in (55), using H 3 in (56), and noticing that $\inf_{\theta\in\Theta}f_{t}(\theta)-\inf_{\theta\in\Theta}\tilde{f}_{t}(\theta)\leqslant\sup_{\theta\in\Theta}f_{t}(\theta)-\tilde{f}_{t}(\theta)$ in (57) leads to:
$\mathfrak{R}_{T}^{\mathrm{d}}$$\leqslant\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle+\xi T$
Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
$\mathfrak{R}_{T}^{\mathrm{d}}$$\leqslant\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}+\xi T$
Observing that $\lVert\bar{g}_{t}(X_{t})\rVert_{\infty}\leqslant\lVert\bar{g}_{t}(X_{t})\rVert_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H 1-(ii), and taking the expectation over $X_{1},\ldots,X_{T}$ and $u_{1},\ldots,u_{T}$ then gives:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ $\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}}\right]+2D_{\mathcal{Z}}\sum_{t\in[T]}\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T\;$ (58)
We now proceed to bound the two first terms in (58). First, notice that by Lemma 1:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]$ $\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$ (59)
Second, noticing that $\tilde{f}_{t}$ is $K_{t}-$Lipschitz with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$ almost surely by Lemma 2, we have by Lemma 5:
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}^{\mathrm{d}}}\right]\leqslant\frac{1}{2\eta_{T}}(D_{\Theta}^{2}+D_{\Theta}P_{T})+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}^{2}G^{2}\eta_{t}}{32\alpha_{t}^{2}}\;.$ (60)

### v2

@@ G.4 Proof of Proposition˜2
For the sake of conciseness we use the notation $\mathbb{E}_{t}$ to denote $\mathbb{E}\mathopen{}\mathclose{{\left[\cdot\mid\mathcal{H}_{t-1}}}\right]$ In what follows, we consider, for any $\theta$, the intermediary regret: $\mathrm{R}_{T}^{s}(\theta):=\sum_{t=1}^{T}F_{t}(\theta_{t})-F_{t}(\theta)$, where $\quad F_{t}:\theta\mapsto\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)\mid\mathcal{H}_{t-1}}}\right]$. We have, for any $\theta$, the following decomposition of the static regret
$\mathrm{R}_{T}^{s}(\theta)$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]$ (49)
@@
Then, taking the $\sup$ over $\Theta$ in Equations˜49 and 51, and defining
@@
in Equation˜50 yields for any $\theta$:
$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$
[…]
$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]$
Since for any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{2}\leqslant D_{\mathcal{Z}}$ by H˜1-(ii),
$\mathrm{R}_{T}^{s}(\theta)$ $\leqslant\mathrm{Reg}^{\mathrm{S}}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\mathbb{E}_{t}{\sup_{\theta\in\Theta}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}\;$
First, notice that, by Lemma˜2, and because for any $t,\alpha_{t}=\alpha$ we have for all $t$:
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}}\right\|_{1}\mid\mathcal{H}_{t-1}}}\right]\leqslant\alpha\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right)\;.$ (52)
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]+2D_{\mathcal{Z}}\alpha T\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right)\;.$ (53)
To bound the first term in the right-hand-side of Equation˜53. by the definition of conditional expectation, we have:
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]$ $=\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})}}\right\rangle-\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}^{\mathrm{S}}_{T}}}\right]$ $\leqslant\mathbb{E}_{X_{1}\cdots X_{T}}\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}}\right].$
@@
One recognises (up to a factor $T$) the left-hand side of Proposition˜4 on the loss sequence $(\tilde{f}_{t})_{t\in[T]}$. Furthermore, we can use this proposition as, given our choice of $\mathcal{R}$, for any $t\in[T]$, $\tilde{f}_{t}$ is $L$-Lipschitz with $L=5D_{\mathcal{Z}}G(4\alpha)^{-1}$ almost surely by Lemma˜1 (with $\chi=0)$. We then have:
@@
$\mathbb{E}_{\sigma_{1},\cdots,\sigma_{T}}\mathopen{}\mathclose{{\left[\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\inf_{\theta\in\Theta}\frac{1}{T}\sum_{t\in[T]}\tilde{f}_{t}(\theta)}}\right]\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi}}\right)\;.$ (54)
Dividing Equation˜53 by $T$, and plugging Equation˜54 gives: for all $\theta\in\Theta$:
@@
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]$ $\leqslant\mathcal{O}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi}}\right)$
$+2D_{\mathcal{Z}}\alpha\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha}}}\right)+(1-\beta)\ln^{2}(\alpha\ln(d))}}\right)}}\right),$
$=\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha\ln(d)}}\right).$
Finally remark that, by the definition of the conditional expectation (thus of $F_{t}$), and because $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable, we have for any $\theta$:
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{R}_{T}^{s}(\theta)}}\right]$ $=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}f_{t}(\theta_{t})-f_{t}(\theta)}}\right]$
$=\mathbb{E}\mathopen{}\mathclose{{\left[\sum_{t=1}^{T}f_{t}(\theta_{t})}}\right]-\mathbb{E}\mathopen{}\mathclose{{\left[f_{t}(\theta)}}\right]$
Then taking the infimum over $\theta$ yields:
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{s}_{T}}}\right]$ $\leqslant\tilde{\mathcal{O}}\mathopen{}\mathclose{{\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{mD}{\eta T}+\xi+\alpha\ln(d)}}\right).$
This concludes the proof. The second equation consists in simply plugging the proposed value of $\eta,\alpha$ in this bound. ∎
G.5 Proof of Proposition˜3
See 3
Proof.
Observe that the dynamic regret can be decomposed as follows:
$\mathfrak{R}^{d}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}}\right\rangle}}\right]+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$ (55)
First, we remark that for any $t$, given $\vartheta_{t}=\mathbf{O}(\tilde{f}_{t})$, we control (56) by Jensen:
$\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]-\inf_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta)}}\right\rangle}}\right]$ $\leqslant\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\tilde{f}_{t}(\vartheta_{t})-\inf_{\theta\in\Theta}\tilde{f}_{t}(\theta)}}\right]$
$\leqslant\xi.$
Then, taking the $\sup$ over $\Theta$ for each summand of the first sum of (55) (which is valid as $\theta_{t}$ is $\mathcal{F}_{t-1}$-measurable), defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}}\right\rangle}}\right]$, and noticing for (57) that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]$ leads to:
$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\rangle}}\right]+\xi T$
[…]
$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]+\xi T$
Since for any $t\in[T]$, $\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{\left\|\bar{g}_{t}(X_{t})}}\right\|_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H˜1-(ii),
$\mathfrak{R}^{d}_{T}$ $\leqslant\mathrm{Reg}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{\left[\mathopen{}\mathclose{{\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}}\right\|_{1}}}\right]+\xi T\;$ (58)
First, notice that, by Lemma˜2, for all $t$:
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{\left[\mathrm{Reg}_{T}}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{\left(\ln\mathopen{}\mathclose{{\left(\frac{2}{\alpha_{t}}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}}\right)}}\right)+\xi T\;.$ (60)
To bound the first term in the right-hand side of Equation˜60. Remark that, by the definition of conditional expectation:
One recognises the definition of $\tilde{\mathfrak{R}^{d}_{T}}$ of Lemma˜8. Thus, by this lemma, we know that for any $t\in[T]$, because $\tilde{f}_{t}$ is $K_{t}$-Lipschitz with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$ almost surely by Lemma˜1, we have: